OpenStudy (anonymous):

The radian measure of an angle that is 25º 47' 31'' is _____. A. .44 radians B. 25.79 radians C. 8.21 radians D. .45 radians

jimthompson5910 (jim_thompson5910):

25º 47' 31'' = 25+47/60 + 31/3600 = 25.7919444444444 Now multiply that last figure by (pi/180) to convert to radians 25.7919444444444*(pi/180) = 0.45015435104701 So 25º 47' 31'' is approximately 0.45 radians.